OpenStudy (anonymous):

How do I find the slope of two-variable linear equation? I'm studying level curves from the stewart text. It says that the level curves of the equation f(x,y) =6-3x-2y are a family of lines with slope -3/2. How?

OpenStudy (amistre64):

how do we measure slope?

OpenStudy (anonymous):

Rise/run? Derivative? Tangent line? :/

OpenStudy (amistre64):



OpenStudy (amistre64):

yeah now what you have is a plane given to you, the 3d version of a line z = 6 - 3x - 2y or 3x + 2y +z = 6, has a normal to the plane defined by the coeeficients (3,2,1) does this make sense?

OpenStudy (perl):

level curves are defined as f(x,y) = k , therefore 6 - 3x - 2y = k

OpenStudy (perl):

that is an equation of a line, solve for y

OpenStudy (perl):

lines* plural, because k is a parameter

OpenStudy (perl):

$$ \Large{ f(x,y) = 6-3x-2y\\ \text{Level curves are defined as } f(x,y) = k \\ \implies \text {by substitution} \\ 6-3x-2y = k } $$

OpenStudy (anonymous):

I see the light now - thanks to both for your perspectives